OpenStudy (anonymous):

approximate these coordinates to the nearest hundreth x^2+y=169 x=12

OpenStudy (anonymous):

x^2+y=169,x=12 x = 12, y = 25

OpenStudy (anonymous):

can u explain how u got that ?

OpenStudy (anonymous):

(12)^2+y = 169 144+y = 169 y = 169-144 y = 25 so,x =12, y=25

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

always welcome